the direction ratio of a line are 2,3,7 then its direction cosiines are

Grade:12

1 Answers

Arun
6 years ago
Dear Mohit
 
We are given with direction ratios (2,3,7)
Its mod = (4+9+49)1/2 = 621/2
Hence direction cosines are
(2/621/2,  3/621/2  , 7/621/2 )
 
Regards
Arun(askIITians forum expert)